OpenStudy (anonymous):

Solve the equation 2a - b = ac + 3b for a.

OpenStudy (anonymous):

get all of the terms with a in them on one side 2a-b-ac=3b

OpenStudy (anonymous):

then 2a-ac=3b+b

OpenStudy (anonymous):



OpenStudy (anonymous):

then factor out the a

OpenStudy (anonymous):

2a - b = ac + 3b get all the (a) variables at one side and you will get 2a-ac=4b take out the common on the right side i.e. (a) a(2-c)=4b a=4b/(2-c)

OpenStudy (anonymous):

a(2-c)=4b--> divide by (2-c) -->(2-c) a=4b/(2-c